mhchen:

Find a sequence of nested open intervals such that A1 is a superset of A2 A2 is a superset of A3 A3 is a superset of A4... Basically, the size of each set is like: A_1 > A_2 > A_3 > A_4... and the union of all of these subsets A_n, as n goes to infinity, is a finite set.

mhchen:

Failed Attempt 1: A_n = (5 - 1/n , 5 + 1/n) A_1 = (5 - 1/1, 6 - 1/1) A_2 = (5 - 1/2, 5 + 1/2) A_3 = (5 - 1/3, 5 + 1/3) A_inf = (5,5) = empty set.

mhchen:

Failed Attempt 2: A_n = (5 - 1/n , 6 + 1/n) A_1 = (5 - 1/1, 6 - 1/1) A_2 = (5 - 1/2, 6 + 1/2) A_3 = (5 - 1/3, 6 + 1/3) A_inf = (5,6) = infinite set

mhchen:

I have concluded that this is impossible. The union of infinite number of sets can only be finite if it converges to a single digit, but since it's an open-interval, the single-digit will not be included in the set. So it will be empty.